# From MRI to SARI: Parametric survey of non-axisymmetric accretion disk instabilities

### v1

@@ Appendix A Frieman-Rotenberg formalism
that contains the static MHD force operator $\mathbf{F}$ as defined in Goedbloed et al. (2019) with additional contributions from the background flow, the equation of motion can be written in terms of $\boldsymbol{\xi}$:
@@
The operator $U=-i\rho\mathbf{v_{0}}\cdot\nabla$ contains the gradient along the background flow and produces a Doppler shift in the spectral equation. Note that Eq. (A4) is quadratic in the eigenvalue $\omega$. However, contrary to popular belief (Ebrahimi & Pharr, 2022), the operators $\mathbf{G}$ and $U$ themselves are self-adjoint (see Chapter 12 of Goedbloed et al. (2019) for a proof), making the difficulty of the spectral problem not one of non-self-adjointness but of a non-linear eigenvalue equation with non-orthogonal eigenfunctions (Goedbloed & Keppens, 2022). In the Spectral Web method, the notion of a complementary energy is used to define exact modes and quasi-modes. This quantity is calculated from the total energy of an open system (Goedbloed et al., 2019):
@@
where $W$ is the total energy, $W^{p}$ the potential energy and the complementary energy $W_{\text{com}}$. The former is real and calculated from the symmetric part of the integrand, whereas the latter can be complex and represents a surface integral over the boundaries. For normal modes, the complementary energy is exactly zero, and so has vanishing real and imaginary parts. These two conditions can be used to construct two paths in the spectral plane on which eigenvalue solutions must lie. Physically, the complementary energy represents additional energy required to bring a mode into resonance with the time dependence $\exp(-i\omega t)$. True normal modes are natural resonances and hence have vanishing $W_{\text{com}}$. In the Spectral Web method, non-zero $W_{\text{com}}$ implies a discontinuity in the total pressure perurbation where the left and right solution obtained from shooting from the boundaries should match. For quasi-continuum SARIs, this jump is so small that the modes can almost be regarded as normal modes.
The vector equation (A4) in $\boldsymbol{\xi}$ can be reduced to a scalar ODE by replacing gradients in the azimuthal and vertical direction by their Fourier-transformed representations, reducing the equation to one for the Fourier coefficients $\hat{\boldsymbol{\xi}}$. The second and third component of $\hat{\boldsymbol{\xi}}$ can then be written in terms of the first component $\xi_{r}$. These expressions can then be used to write the equation of motion in terms of $\chi=r\xi_{r}$ alone as a second-order ODE,
@@
This generalised Hain-Lüst equation was first obtained in the context of accretion disks by Keppens et al. (2002) in a different form, and rewritten by Blokland et al. (2005) in the form above. It contains function coefficients which are in general quite complicated expressions involving equilibrium quantities, the wavenumbers $m$ and $k$ and the frequency $\omega$. They can be found in Goedbloed et al. (2019) and Goedbloed & Keppens (2022), but we will repeat the most important definitions here.
@@
The singular solutions of Eq. (A6) are produced by coefficients $N$ and $D$. Actually, the latter function only leads to apparent singularities, which becomes clear after rewriting Eq. (A6) as a system of first-order ODEs (Blokland et al., 2005). The genuine singularities, resulting from the zeroes of $N$, can be obtained explicitly,
@@
The four singular ranges are dubbed forward and backward Alfvén and slow continuum, respectively, as the frequencies lie in a continuum band of Alfvén frequencies $\omega_{A}^{2}$ and slow frequencies $\omega_{S}^{2}=\omega_{A}^{2}\ \gamma p/(\gamma p+B^{2})$ shifted by the Doppler range $\Omega_{0}$. The eigenfunctions corresponding to these continuum frequencies are non-square integrable functions with a singularity at the location of resonance with the continuum, i.e. where $\Re(\tilde{\omega})=\omega_{A,S}(r_{\text{res}})$.
@@ Appendix B Dispersion relation
In the global formulation above, the eigenfunctions have a radial dependence that obeys the second-order ODE (A6). In the numerical survey in the main body of this work, we found that eigenfunctions of MRI, discrete SARI and quasi-continuum modes clearly differ, with each having their own typical radial variation and (in)sensitivity to boundary conditions. In the local limit, however, a dispersion relation can be derived assuming a purely sinusoidal radial dependence. This Section connects some recent results on accretion disk dispersion relations and non-axisymmetric instabilities.
@@
the functions $\tilde{\chi}$ and $q$ can be determined from the real and imaginary parts of the equation and using the WKB assumption (Blokland et al., 2005):
@@
where the coefficients can be found in Blokland et al. (2005). It describes the relation between $\tilde{\omega}$, the wavenumbers $(q,m,k)$ and the equilibrium quantities at every location $r\in[r_{1},r_{2}]$. This dispersion relation contains various local dispersion relations obtained for specific approximate cases, including the original standard weak-field MRI dispersion relation by Balbus & Hawley (1991) as shown in Blokland et al. (2005), and fourth and sixth-order relations by Kim & Ostriker (2000) and by Balbus & Hawley (1998), respectively, as demonstrated in Keppens et al. (2002). The sixth-order dispersion relation (B4) describes any general cylindrical disk equilibrium for any $(m,k)$-combination. Its six solutions consist of four oscillating modes (the Alfvén and fast modes) and a conjugate pair of unstable modes (MRI- or SARI-like modes). It was originally applied to the overstability of the $m=0$ mode, which is well described by this WKB approach.
@@ B.1 Recent dispersion relations
Das et al. (2018) derived a general fourth-order dispersion relation from the linearised MHD equations, incorporating a suprathermal poloidal field, curvature, compressibility and non-axisymmetric modes (their Eq. (40)). It is indeed contained in the description by Blokland et al. (2005) and can be obtained by applying the following approximations to Eq.(B4):
@@
  • neglecting the equilibrium variations of the thermodynamic variables, i.e. set $\rho^{\prime}=p^{\prime}=0$.
These final two assumptions allow us to simplify the deviation from Keplerian force balance $\Lambda$ as,
@@
and after rearranging terms, their dispersion relation (40) is obtained. Note that the first condition implies the WKB assumption $q^{2}r^{2}~{}\gg~{}1$. It ensures that curvature terms from radial derivatives of the perturbations are ignored with respect to the background. Das et al. (2018) compared the solutions of this dispersion relation to solutions of the global eigenvalue problem in the limit of a dominant suprathermal toroidal field. Their main focus was to extend and confirm local predictions by Pessah & Psaltis (2005), who found that the MRI is suppressed entirely over all wavenumbers for strong enough $v_{A\theta 1}$. The global study does not find a complete stabilisation, but rather very small growth rates for $k<k_{\text{cut-off}}$. However, two instabilities appear for higher toroidal field that are unstable slow and hybrid slow-Alfvén modes, and have overlapping ranges of unstable wavelengths $k$. These modes are also predicted by the dispersion relation, but their cut-off wavenumbers do not overlap. Our Case (b) for the growth rate variation with $\beta$ (Sect. 5.3) is somewhat similar to their set-up, but we do not observe a stabilisation of the MRI at a certain field strength, the reason being that while $\beta$ is varied, $v_{A1}$ remains constant and hence it is the sound speed that varies. Note that in all cases these authors assume $m=0$, so there is of course no occurrence of a quasi-continuum.
@@
Recently, Ebrahimi & Pharr (2022) considered non-axisymmetric instabilities in a cylindrical equilibrium flow relevant for MRI laboratory experiments. The authors similarly arrive at a local dispersion relation including curvature and non-axisymmetric effects. There are, however, several notable differences with the analysis presented in Blokland et al. (2005) and Das et al. (2018).
@@
Perhaps the most distinctive difference is that the adopted stationary equilibrium significantly deviates: in Ebrahimi & Pharr (2022), balance against the centrifugal force is provided solely by the plasma pressure, since the flow is Keplerian with a currentless (potential) magnetic field, no gravity and constant radial density. The pressure – and hence disk temperature – is hence implicitly forced to be radially increasing through Eq. (2), in contrast to the negative power laws used in Blokland et al. (2005), Das et al. (2018) and Goedbloed & Keppens (2022). This setup, where a rotating plasma column is pressure-balanced, may well be more relevant for laboratory experiments than for physical accretion disks. However, their fourth-order dispersion relation Eq. (7) contains various terms that can not directly be reconciled with any of the aforementioned works. Their dispersion relation is derived from linearised momentum and induction equations, but incompressibility is implicitly assumed and eigenfunctions $\rho_{1}$, $p_{1}$, $v_{z1}$ and $B_{z1}$ have been eliminated. The authors argue for introducing a curvature term that is not included in previous works, replacing the radial wavenumber $q$ by $\bar{q}:=q+1/r$. However, inclusion of such a ‘curvature’ term $1/r$, which relates to the radial derivative in cylindrical coordinates, introduces imaginary terms $\sim iq$ that are not present in other published dispersion relations. In a true WKB approximation, the wavenumber $q$ is not modified because of the assumption $qr\gg 1$, and curvature from the cylindrical geometry is only included in the dispersion relation through the azimuthal components $v_{\theta}/r$ and $B_{\theta}/r$. With this approximation, the factors $iq$ are still not eliminated from the dispersion relation. In the derivation of Das et al. (2018), the following two assumptions prevent these factors from appearing:
@@
The first assumption results from neglecting $v_{z1}$ and $B_{z1}$ in the vertical momentum equation (justified by Eq. (2.22) of Balbus & Hawley (1991)), while the second assumption follows from the ordering $q^{2}r^{2}\sim k^{2}r^{2}\gg m^{2}$ applied to $\nabla\cdot\mathbf{B_{1}}=0$. With mutual consistency of the Blokland et al. (2005) and Das et al. (2018) results, we hence conclude that the inclusion of terms $\sim iq$ is not justified. Das et al. (2018) came to a similar conclusion when comparing their dispersion relation to that of Pessah & Psaltis (2005).

### v2

@@ Appendix A Frieman-Rotenberg formalism
which contains the static MHD force operator $\mathbf{F}$ as defined in Goedbloed et al. (2019) with additional contributions from the background flow, the equation of motion can be written in terms of $\boldsymbol{\xi}$:
@@
The operator $U=-i\rho\mathbf{v_{0}}\cdot\nabla$ contains the gradient along the background flow and produces a Doppler shift in the spectral equation. Note that Eq. (A4) is quadratic in the eigenvalue $\omega$. However, contrary to popular belief (Ebrahimi & Pharr, 2022), the operators $\mathbf{G}$ and $U$ themselves are self-adjoint (see Chapter 12 of Goedbloed et al. (2019) for a proof), making the difficulty of the spectral problem not one of non-self-adjointness but one of a non-linear eigenvalue equation with non-orthogonal eigenfunctions (Goedbloed & Keppens, 2022). In the Spectral Web method, the notion of a complementary energy is used to define exact modes and quasi-modes. This quantity is calculated from the total energy of an open system (Goedbloed et al., 2019):
@@
where $W$ is the total energy, $W^{p}$ is the potential energy and $W_{\text{com}}$ is the complementary energy . The former is real and calculated from the symmetric part of the integrand, whereas $W_{\text{com}}$ can be complex and represents a surface integral over the boundaries. For normal modes, the complementary energy is exactly zero, and so has vanishing real and imaginary parts. These two conditions can be used to construct two paths in the spectral plane on which eigenvalue solutions must lie. Physically, the complementary energy represents the additional energy required to bring a mode into resonance with the time dependence $\exp(-i\omega t)$. True normal modes are natural resonances and hence have vanishing $W_{\text{com}}$. In the Spectral Web method, non-zero $W_{\text{com}}$ implies a discontinuity in the total pressure perturbation where the left and right solutions obtained from shooting from the boundaries should match. For quasi-continuum SARIs, this jump is so small that the modes can almost be regarded as normal modes.
The vector equation (A4) in $\boldsymbol{\xi}$ can be reduced to a scalar ODE by replacing gradients in the azimuthal and vertical directions by their Fourier-transformed representations, reducing the equation to one for the Fourier coefficients $\hat{\boldsymbol{\xi}}$. The second and third components of $\hat{\boldsymbol{\xi}}$ can then be written in terms of the first component $\xi_{r}$. These expressions can then be used to write the equation of motion in terms of $\chi=r\xi_{r}$ alone as a second-order ODE:
@@
This generalised Hain-Lüst equation was first obtained in the context of accretion disks by Keppens et al. (2002) in a different form, and was rewritten by Blokland et al. (2005) in the form above. It contains function coefficients that are in general quite complicated expressions involving equilibrium quantities, the wavenumbers $m$ and $k$ and the frequency $\omega$. They can be found in Goedbloed et al. (2019) and Goedbloed & Keppens (2022), but we will repeat the most important definitions here.
@@
The singular solutions of Eq. (A6) are produced by the coefficients $N$ and $D$. Actually, the latter function only leads to apparent singularities, which becomes clear after rewriting Eq. (A6) as a system of first-order ODEs (Blokland et al., 2005). The genuine singularities, resulting from the zeroes of $N$, can be obtained explicitly:
@@
The four singular ranges are dubbed forward and backward Alfvén and slow continua, as the frequencies lie in a continuum band of Alfvén frequencies $\omega_{A}^{2}$ and slow frequencies $\omega_{S}^{2}=\omega_{A}^{2}\ \gamma p/(\gamma p+B^{2})$ shifted by the Doppler range $\Omega_{0}$. The eigenfunctions corresponding to these continuum frequencies are non-square integrable functions with a singularity at the location of resonance with the continuum, i.e. where $\Re(\tilde{\omega})=\omega_{A,S}(r_{\text{res}})$.
@@ Appendix B Dispersion relation
In the global formulation above, the eigenfunctions have a radial dependence that obeys the second-order ODE (A6). In the numerical survey in the main body of this work, we found that the eigenfunctions of MRI, discrete SARI, and quasi-continuum modes clearly differ, each having its own typical radial variation and (in)sensitivity to boundary conditions. In the local limit, however, a dispersion relation can be derived assuming a purely sinusoidal radial dependence. This Section connects some recent results on accretion disk dispersion relations and non-axisymmetric instabilities.
@@
the functions $\tilde{\chi}$ and $q$ can be determined from the real and imaginary parts of the equation and by using the WKB assumption (Blokland et al., 2005):
@@
where the coefficients can be found in Blokland et al. (2005). It describes the relation between $\tilde{\omega}$, the wavenumbers $(q,m,k)$, and the equilibrium quantities at every location $r\in[r_{1},r_{2}]$. This dispersion relation contains various local dispersion relations obtained for specific approximate cases, including the original standard weak-field MRI dispersion relation by Balbus & Hawley (1991) as shown in Blokland et al. (2005), and fourth- and sixth-order relations by Kim & Ostriker (2000) and by Balbus & Hawley (1998), respectively, as demonstrated in Keppens et al. (2002). The sixth-order dispersion relation (B4) describes any general cylindrical disk equilibrium for any $(m,k)$-combination. Its six solutions consist of four oscillating modes (the Alfvén and fast modes) and a conjugate pair of unstable modes (MRI- or SARI-like modes). It was originally applied to the overstability of the $m=0$ mode, which is well described by this WKB approach.
@@ B.1 Recent dispersion relations
Das et al. (2018) derived a general fourth-order dispersion relation from the linearised MHD equations, incorporating a suprathermal poloidal field, curvature, compressibility, and non-axisymmetric modes (their Eq. (40)). It is indeed contained in the description by Blokland et al. (2005) and can be obtained by applying the following approximations to Eq.(B4):
@@
  • neglecting the equilibrium variations of the thermodynamic variables, i.e. setting $\rho^{\prime}=p^{\prime}=0$.
These final two assumptions allow us to simplify the deviation from Keplerian force balance $\Lambda$ as
@@
and after rearranging terms, their dispersion relation (40) is obtained. Note that the first condition implies the WKB assumption $q^{2}r^{2}~{}\gg~{}1$. It ensures that curvature terms from radial derivatives of the perturbations are ignored with respect to the background. Das et al. (2018) compared the solutions of this dispersion relation to solutions of the global eigenvalue problem in the limit of a dominant suprathermal toroidal field. Their main focus was to extend and confirm local predictions by Pessah & Psaltis (2005), who found that the MRI is suppressed entirely over all wavenumbers for strong enough $v_{A\theta 1}$. The global study does not find a complete stabilisation, but rather very low growth rates for $k<k_{\text{cut-off}}$. However, two instabilities appear for higher toroidal fields that are unstable slow and hybrid slow-Alfvén modes, and have overlapping ranges of unstable wavelengths $k$. These modes are also predicted by the dispersion relation, but their cut-off wavenumbers do not overlap. Our Case (b) for the growth rate variation with $\beta$ (Sect. 5.3) is somewhat similar to their set-up, but we do not observe a stabilisation of the MRI at a certain field strength, the reason being that while $\beta$ is varied, $v_{A1}$ remains constant and hence it is the sound speed that varies. Note that in all cases these authors assume $m=0$, so there is of course no occurrence of a quasi-continuum.
@@
Recently, Ebrahimi & Pharr (2022) considered non-axisymmetric instabilities in a cylindrical equilibrium flow relevant for MRI laboratory experiments. The authors similarly arrived at a local dispersion relation including curvature and non-axisymmetric effects. There are, however, several notable differences from the analysis presented in Blokland et al. (2005) and Das et al. (2018).
@@
Perhaps the most distinct difference is that the adopted stationary equilibrium significantly deviates: in Ebrahimi & Pharr (2022), balance against the centrifugal force is provided solely by the plasma pressure, since the flow is Keplerian with a currentless (potential) magnetic field, no gravity, and a constant radial density. The pressure – and hence the disk temperature – is hence implicitly forced to be radially increasing through Eq. (2), in contrast to the negative power laws used in Blokland et al. (2005), Das et al. (2018) and Goedbloed & Keppens (2022). This setup, where a rotating plasma column is pressure-balanced, may well be more relevant for laboratory experiments than for physical accretion disks. However, their fourth-order dispersion relation Eq. (7) contains various terms that cannot directly be reconciled with any of the aforementioned works. Their dispersion relation is derived from linearised momentum and induction equations, but incompressibility is implicitly assumed and eigenfunctions $\rho_{1}$, $p_{1}$, $v_{z1}$ and $B_{z1}$ have been eliminated. The authors argue for introducing a curvature term that is not included in previous works, replacing the radial wavenumber $q$ by $\bar{q}:=q+1/r$. However, the inclusion of such a ‘curvature’ term $1/r$, which relates to the radial derivative in cylindrical coordinates, introduces imaginary terms $\sim iq$ that are not present in other published dispersion relations. In a true WKB approximation, the wavenumber $q$ is not modified because of the assumption $qr\gg 1$, and curvature from the cylindrical geometry is only included in the dispersion relation through the azimuthal components $v_{\theta}/r$ and $B_{\theta}/r$. With this approximation, the factors $iq$ are still not eliminated from the dispersion relation. In the derivation of Das et al. (2018), the following two assumptions prevent these factors from appearing:
@@
The first assumption results from neglecting $v_{z1}$ and $B_{z1}$ in the vertical momentum equation (justified by Eq. (2.22) of Balbus & Hawley (1991)), while the second assumption follows from the ordering $q^{2}r^{2}\sim k^{2}r^{2}\gg m^{2}$ applied to $\nabla\cdot\mathbf{B_{1}}=0$. With the mutual consistency of the Blokland et al. (2005) and Das et al. (2018) results, we conclude that the inclusion of terms $\sim iq$ is not justified. Das et al. (2018) came to a similar conclusion when comparing their dispersion relation to that of Pessah & Psaltis (2005).